OpenStudy (clamin):

PLEASE HELP!! What number must you add to complete the square? x^2 +10x =37 a. 10 b. 20 c. 25 d. 5

OpenStudy (isaidavila):

(10/2)^2

OpenStudy (clamin):

c @isaidavila

OpenStudy (isaidavila):

Yes

OpenStudy (clamin):

@isaidavila can you help me with this one??

OpenStudy (abdullahm):

Hint: \(\bf a^2 + 2ab + b^2 = (a+b)^2\\a^2-2ab+b^2 = (a-b)^2\) So we have 3x in place of 2ab, and we know that a = x. So we have: 3x = 2xb 3 = 2b what is 'b'? and then whatever you get, we need to square that because we need to find \(\bf b^2\)

OpenStudy (clamin):

-3

OpenStudy (abdullahm):

What did you find as "-3" Hint: 3 = 2b Solve for b, and then square it to find \(\bf b^2\)

OpenStudy (clamin):

-3/2

OpenStudy (clamin):

is it right?? @AbdullahM

OpenStudy (abdullahm):

and then square that hint: \(\bf\Large\left(\frac{a}{b}\right)^2 =\frac{a^2}{b^2}\)

OpenStudy (clamin):

9/4 d.

OpenStudy (abdullahm):

Correct!